Suppose 20% of grade-school students develop influenza. In a sample of 100 grade-school students, what is the probability that 30 or more of them will develop influenza?
Answer #1

mean = np = 100 * 20% = 20

standard error = sqrt(100*20%*80%) = 4

P(X >=30) = P( Z> 30-20/4)

= P(Z > 2.5)

= 0.0062